A proton accelerates from rest in a uniform electric field of 660 N/C. At one later moment, its speed is 1.40 Mm/s (nonrelativistic because v is much less than the speed of light).

(a) Find the acceleration of the proton.
m/s2

(b) Over what time interval does the proton reach this speed?
s

(c) How far does it move in this time interval?
m

(d) What is its kinetic energy at the end of this interval?
J
Answer #1
Part a

The acceleration is 6.32×1010m/s26.32 \times {10^{10}}{\rm{ m/}}{{\rm{s}}^2}.

Part b

The time taken by proton to reach its final speed is 2.2×105s2.2 \times {10^{ - 5}}{\rm{ s}}.

Part c

The distance moved by the proton in time t is 15.3 m.

Part d

The kinetic energy of proton is 1.63×1015J1.63 \times {10^{ - 15}}{\rm{ J}}.